### v3

# Computation of the deformation of rhombi-slit kirigami

Abstract
Kirigami are part of the larger class of mechanical metamaterials, which exhibit exotic properties. This article focuses on rhombi-slits, which is a specific type of kirigami. A nonlinear kinematic model was previously proposed as a second order divergence-form PDE with a possibly degenerate, and sign-changing coefficient matrix. We first propose to study the existence of solutions to a regularization of this equation by using the limiting absorption principle. Then, we propose a finite element method with complex polynomials to approximate the solutions to the nonlinear equation. Finally, simulations are compared with experimental results.
Keywords: Kirigami, Degenerate PDE, Sign-changing PDE, Limiting absorption principle.
AMS Subject Classification: 35M12, 65N12, 65N30 and 35Q74.
1 Introduction
Kirigami is a variation of origami, the Japanese art of paper folding. In kirigami, the paper is cut as well as being folded. Origami and kirigami have been studied as concrete examples of mechanical metamaterials [24, 2, 31, 28]. Mechanical metamaterials are solids with unusual mechanical properties that are generated by inner mechanisms of the system. These materials have found applications for generating soft robots [23] or in aerospace engineering [18, 22]. The mechanisms giving metamaterials their unusual properties are discrete by nature, and have been modeled as such [29, 11]. Recently, important efforts have been devoted to proposing a continuous description of origami and kirigami-based metamaterials [19, 14, 20, 30]. The modeling process results in homogenized PDEs where the characteristic size of the mechanism is considered negligible with respects to the size of the whole structure. Some first results regarding the existence of solutions to these nonlinear PDEs, and their numerical approximation have been achieved in [17, 16, 30].
This paper focuses on rotating square patterns as presented in [30]. Figure 1 shows one such kirigami pattern. In this entire article, deformations remain planar. $\xi_{\mathrm{in}}$ is defined as the slit opening in the undeformed configuration. $\xi$ is the opening of the slit, and $\gamma$ is the local rotation of a panel, as represented in Figure 1.
$2\xi_{\mathrm{in}}$
$\gamma$$2(\xi+\xi_{\mathrm{in}})$
Figure 1: Notations.
$y_{\mathrm{eff}}$ is the effective deformation, which tracks the cell-averaged panel motions. Using a coarse-graining technique, [30] has proposed that
$\nabla y_{\mathrm{eff}}=R(\gamma)A_{\mathrm{eff}}(\xi),$ (1)
where $R(\gamma)$ is the canonical $2\times 2$ rotation matrix parametrized by the angle $\gamma$, and $A_{\mathrm{eff}}(\xi)$ is called the shape tensor. It is shown in [30] that taking the curl of (1) gives
$\nabla\gamma=\Gamma(\xi)\nabla\xi,$ (2)
where $\Gamma(\xi)$ is a $2\times 2$ matrix, which depends on $A_{\mathrm{eff}}(\xi)$. Taking the curl of (2), which is equivalent to taking the divergence after a rotation by $\frac{\pi}{2}$, one gets
$-\mathrm{div}\left(R\left(\frac{\pi}{2}\right)\Gamma(\xi)\nabla\xi\right)=0.$ (3)
The main goal of this paper is to compute approximate solutions to (3) supplemented with appropriate Dirichlet and Neumann boundary conditions. In [30], the authors compute solutions of (3) only in two two specific cases: $\alpha=-\beta$ and $\beta=0$ in the notation presented below. In the follow-up work [31], the authors give an energy and compute its minimizers but they do not concurrently solve the kinematic constraint. The present work proposes a way to approximate the solutions of (3) for any value of these parameters.
Let $B(\xi):=R\left(\frac{\pi}{2}\right)\Gamma(\xi)$. Studying the existence of solutions to (3) presents two main difficulties. The first is that $B(\xi)$ can degenerate for certain values of $\xi$. The second is that the sign of $\mathrm{det}(B(\xi))$ can change, thus locally changing the type of (3) from an elliptic to a hyperbolic PDE. A similar problem is encountered with Tricomi and Keldysh equations, see [6] for instance. Regarding the issue of degeneracy, Muckenhopt weights [27] have been used to study the existence of solutions to degenerate elliptic equations, see [9], for instance. Muckenhopt weights can be used, for instance, to handle degeneracies localized at a point. However, as in [21], the degeneracy we encounter is localized on a curve, and thus Muckenhopt weights cannot be used. Regarding the issue of sign-change, the concept of $\mathtt{T}$-coercivity [3] has been used to study the existence of solutions for some linear equations [3, 7]. Unfortunately, this concept does not seem to be applicable here. Instead, we follow [21, 8] and use a limiting absorption principle to approximate solutions of (3).
Several numerical methods have been proposed to approximate solutions of sign-changing problems. [7] proposed methods based on first-order Lagrange polynomials and $\mathtt{T}_{h}$-coercive meshes. However, these cannot be used in the present case as the position of the curve where (3) changes type is not known, a priori. More recently, [5] proposed a method based on optimal trasnport. However, it cannot be applied as we deal with a degenerate equation. Therefore, we compute solutions of a problem regularized by adding a complex dissipation, as already performed in [7].
The present article is structured as follows. Section 2 focuses on proving the existence of solutions for a regularization of (3) by using the the limiting absorption principle. Section 3 proposes a numerical method to approximate the solutions of the regularization of (3), supplemented by a convergence proof. Finally, Section 4 compares a few numerical tests to experimental results.
2 Approximate continuous problem
Let $\Omega\subset\mathbb{R}^{2}$ be an open bounded polygonal domain, that can be perfectly fitted by triangular meshes, with a Lipschitz boundary $\partial\Omega$. The boundary is partitioned as $\partial\Omega=\partial\Omega_{D}\cup\partial\Omega_{N}$, where $\partial\Omega_{D}$ is relatively closed in $\partial\Omega$. A Dirichlet boundary $\xi_{D}\in H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})$ is imposed strongly on $\partial\Omega_{D}$ and a Neumann boundary condition $g\in L^{2}(\partial\Omega_{N};\mathbb{R})$ is imposed weakly on $\partial\Omega_{N}$. Let $V:=H^{1}(\Omega;\mathbb{C})$ be a complex Sobolev space equipped with the usual Sobolev norm. We define the solution space as $V_{D}:=\{\zeta\in V|\Re(\zeta)=\xi_{D},\text{ and }\Im(\zeta)=0\text{ on }\partial\Omega_{D}\}$, and $V_{0}:=\{\zeta\in V|\zeta=0\text{ on }\partial\Omega_{D}\}$ as the associated homogeneous space.
2.1 Rhombi-slits
In the following, we focus on the specific case of rhombi slits, see [30]. Let $\alpha\leq 0$, and $\beta\geq 0$ be geometric parameters which depend on $\xi_{in}$ and the geometry of the panel, see [30]. Let $\xi\in L^{2}(\Omega;\mathbb{R})$. We define
$\mu_{1}(\xi)=\cos(\xi)-\alpha\sin(\xi),\quad\mu_{2}(\xi)=\cos(\xi)+\beta\sin(\xi),$
and
$A(\xi)=\mu_{1}(\xi)e_{1}\otimes e_{1}+\mu_{2}(\xi)e_{2}\otimes e_{2}.$
Let $\Gamma(\xi)=\Gamma_{12}(\xi)e_{1}\otimes e_{2}+\Gamma_{12}(\xi)e_{1}\otimes e_{2}$, where $(e_{1},e_{2})$ is the canonical basis of $\mathbb{R}^{2}$, and
$\Gamma_{12}(\xi)=-\frac{\mu_{1}^{\prime}(\xi)}{\mu_{2}(\xi)},\quad\Gamma_{21}(\xi)=\frac{\mu_{2}^{\prime}(\xi)}{\mu_{1}(\xi)}.$
Finally, one has
$B(\xi)=\begin{pmatrix}-\Gamma_{21}(\xi)&0\\ 0&\Gamma_{12}(\xi)\\ \end{pmatrix}.$
We want to find $\xi$ such that
{$&-\mathrm{div}(B(\xi)\nabla\xi)=0\text{ a.e. in }\Omega,\\ &B(\xi)\nabla\xi\cdot n=g\text{ on }\partial\Omega_{N},\\ &\xi=\xi_{D}\text{ on }\partial\Omega_{D},\end{aligned}\right.$ (4)
where $n$ is the exterior normal to $\partial\Omega$. Note that, defined as such, the coefficients of $B(\xi)$ can take infinite values. Let $\xi^{-}\in[-\frac{\pi}{2},0]$, and $\xi^{+}\in[0,\frac{\pi}{2}]$. We define the cut-off coefficients
$\widehat{\Gamma_{21}}(\xi)=\left\{\begin{aligned} &\Gamma_{21}(\xi^{+})\text{ if }\xi\geq\xi^{+},\\ &\Gamma_{21}(\xi^{-})\text{ if }\xi\leq\xi^{-},\\ &\Gamma_{21}(\xi)\text{ otherwise},\end{aligned}\right.\quad\widehat{\Gamma_{12}}(\xi)=\left\{\begin{aligned} &\Gamma_{12}(\xi^{+})\text{ if }\xi\geq\xi^{+},\\ &\Gamma_{12}(\xi^{-})\text{ if }\xi\leq\xi^{-},\\ &\Gamma_{12}(\xi)\text{ otherwise},\end{aligned}\right.$
and
$\hat{B}(\xi)=\begin{pmatrix}-\widehat{\Gamma_{21}}(\xi)&0\\ 0&\widehat{\Gamma_{12}}(\xi)\\ \end{pmatrix}.$
We choose $\xi^{-}$ and $\xi^{+}$ such that there exists $M>0$, independent of $\xi$,
$\|\hat{B}(\xi)\|_{L^{\infty}(\Omega;\mathbb{R})}\leq M.$ (5)
Note that $\hat{B}$ is $K$-Lipschitz continuous, where $K>0$. Of course, $\hat{B}(\xi)$ is consistent with $B(\xi)$ only as long as $\xi^{-}\leq\xi\leq\xi^{+}$ a.e. in $\Omega$. The bilinear form $a(\xi)$ is defined for $\zeta\in V$ and $\tilde{\zeta}\in V_{0}$ by
$a(\xi;\zeta,\tilde{\zeta}):=\int_{\Omega}\hat{B}(\xi)\nabla\zeta\cdot\overline{\nabla\tilde{\zeta}}.$
Let $l$ be the linear form defined for $\tilde{\zeta}\in V_{0}$ by
$l(\tilde{\zeta}):=\int_{\partial\Omega_{N}}g\cdot\overline{\tilde{\zeta}}.$
2.2 Linear problem
Let $\xi\in L^{2}(\Omega;\mathbb{R})$ be fixed. Because $\mathrm{div}(\hat{B}(\xi))$ can change sign, $a(\xi)$ is not coercive over $V_{0}\times V_{0}$. Therefore, one cannot apply a Lax–Milgram Lemma directly. Instead, we follow [7, 8] and use a limiting absorption principle. In consists in adding some dissipation to $a(\xi)$ through a purely imaginary term. Let $\varepsilon>0$ be a regularization parameter. For $\zeta\in V$ and $\tilde{\zeta}\in V_{0}$, we define
$a_{\varepsilon}(\xi;\zeta,\tilde{\zeta}):=\int_{\Omega}(\hat{B}(\xi)+i\varepsilon I)\nabla\zeta\cdot\overline{\nabla\tilde{\zeta}},$
where $i^{2}=-1$.
Lemma 1.
For all $\varepsilon>0$, there exists a unique solution $\zeta_{\varepsilon}\in V_{D}$ to
$a_{\varepsilon}(\xi;\zeta_{\varepsilon},\tilde{\zeta})=l(\tilde{\zeta}),\quad\forall\tilde{\zeta}\in V_{0}.$ (6)
One has
$\|\zeta_{\varepsilon}\|_{V}\leq\left(\frac{M\sqrt{1+C}}{\varepsilon}+1\right)C_{tr}\|\xi_{D}\|_{H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})}+\frac{\sqrt{1+C}}{\varepsilon}\|g\|_{L^{2}(\partial\Omega_{N};\mathbb{R})},$ (7)
where $C,C_{tr}>0$ are constants independent of $\xi$.
Proof.
As $\xi_{D}\in H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})$, there exists $\zeta_{D}\in H^{1}(\Omega;\mathbb{R})$, $\zeta_{D}=\xi_{D}$ on $\partial\Omega_{D}$. We define the linear form $L(\xi)$ for $\tilde{\zeta}\in V_{0}$ by
$L(\xi;\tilde{\zeta})=l(\tilde{\zeta})-\int_{\Omega}\hat{B}(\xi)\nabla\zeta_{D}\cdot\overline{\nabla\tilde{\zeta}}.$
$a_{\varepsilon}(\xi)$ is sesquilinear continuous over $V_{0}\times V_{0}$. Let $\zeta\in V_{0}$, one has
$\Im(a_{\varepsilon}(\xi;\zeta,\zeta))\geq\varepsilon\|\nabla\zeta\|_{L^{2}(\Omega;\mathbb{C})}^{2}.$
One also has
$\Re(a_{\varepsilon}(\xi;\zeta,\zeta))\geq-M\|\nabla\zeta\|_{L^{2}(\Omega;\mathbb{C})}^{2}.$
Let $C>\frac{M}{\varepsilon}$, one has
$\Re((1-iC)a_{\varepsilon}(\xi;\zeta,\zeta))\geq(C\varepsilon-M)\|\nabla\zeta\|_{L^{2}(\Omega;\mathbb{C})}^{2}.$
Therefore, $a_{\varepsilon}(\xi)$ is coercive over $V_{0}\times V_{0}$. As $L(\xi)$ is continuous over $V_{0}$, we apply the Lax–Milgram lemma and get the existence of $\hat{\zeta_{\varepsilon}}\in V_{0}$ such that $\zeta_{\varepsilon}:=\hat{\zeta}_{\varepsilon}+\zeta_{D}$ solves (6). Let us now prove the bound (7). Testing with $\hat{\zeta}_{\varepsilon}$, and taking the imaginary part, one has
$\frac{\varepsilon}{\sqrt{1+C}}\|\hat{\zeta}_{\varepsilon}\|_{V}^{2}$ $\leq\|\bar{B}(\xi)\nabla\zeta_{D}\|_{L^{2}(\Omega;\mathbb{R})}\|\hat{\zeta}_{\varepsilon}\|_{V}+\|g\|_{L^{2}(\partial\Omega_{N};\mathbb{R})}\|\hat{\zeta}_{\varepsilon}\|_{V},$
$\leq MC_{tr}\|\xi_{D}\|_{H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})}\|\hat{\zeta}_{\varepsilon}\|_{V}+\|g\|_{L^{2}(\partial\Omega_{N};\mathbb{R})}\|\hat{\zeta}_{\varepsilon}\|_{V},$
where $C>0$ is the Poincaré constant and $C_{tr}>0$ is the constant from the trace theorem [10, Theorem B.52], from which one deduces (7). ∎
2.3 Nonlinear problem
Using the results from Section 2.2, we prove that there exists a solution to the approximate problem.
Proposition 2.
For $\varepsilon>0$, there exists a fixed point $\xi_{\varepsilon}\in V_{D}$ solution of
$a_{\varepsilon}(\Re(\xi_{\varepsilon});\xi_{\varepsilon},\tilde{\zeta})=l(\tilde{\zeta}),\quad\forall\tilde{\zeta}\in V_{0}.$ (8)
Proof.
Let $\varepsilon>0$ and $T_{\varepsilon}:L^{2}(\Omega;\mathbb{C})\to V$ be the map such that for $\xi\in L^{2}(\Omega;\mathbb{C})$, $T_{\varepsilon}\xi=\zeta_{\varepsilon}\in V_{D}$ is the unique solution of
$a_{\varepsilon}(\Re(\xi_{\varepsilon});\zeta_{\varepsilon},\tilde{\zeta})=l(\tilde{\zeta}),\quad\forall\tilde{\zeta}\in V_{0},$
[…]
Let us now show that $T_{\varepsilon}$ is continuous over $B$ for $\|\cdot\|_{V}$. Let $\left(\xi_{n}\right)_{n\in\mathbb{N}}$ be a sequence of $B$ such that $\xi_{n}\mathop{\longrightarrow}\limits_{n\to+\infty}\xi\in B$ in $V$. Let $\zeta=T_{\varepsilon}\xi$, and $\zeta_{n}=T_{\varepsilon}\xi_{n}$, for $n\in\mathbb{N}$. We want to prove that $\zeta_{n}\mathop{\longrightarrow}\limits_{n\to+\infty}\zeta$ in $V$. $\left(\zeta_{n}\right)_{n}$ is bounded in $V_{D}$, and thus there exists $\hat{\zeta}\in V_{D}$, up to a subsequence, $\zeta_{n}\mathop{\rightharpoonup}\limits_{n\to+\infty}\hat{\zeta}$ weakly in $V$. Using the classical Sobolev injection $V\subset L^{2}(\Omega;\mathbb{C})$, $\zeta_{n}\to\hat{\zeta}$ strongly in $L^{2}(\Omega;\mathbb{C})$, when $n\to+\infty$. Let us show that $\hat{\zeta}$ is a solution of (6). By definition, one has for all $\tilde{\zeta}\in V_{0}$,
$l(\tilde{\zeta})=a_{\varepsilon}(\Re(\xi_{n});\zeta_{n},\tilde{\zeta})\mathop{\longrightarrow}_{n\to\infty}a_{\varepsilon}(\Re(\xi);\hat{\zeta},\tilde{\zeta}),$
since $\hat{B}$ is $K$-Lipschitz, and thus continuous. Therefore, $\hat{\zeta}\in V_{D}$ solves (6), which has for unique solution $\zeta\in V_{D}$, and thus $\hat{\zeta}=\zeta$. Let us show that $\zeta_{n}\to\zeta$ strongly in $V$ when $\varepsilon\to 0$. One has
$i\varepsilon\|\nabla\zeta_{n}-\nabla\zeta\|_{L^{2}(\Omega;\mathbb{C})}^{2}+\int_{\Omega}\hat{B}(\Re(\xi_{n}))\nabla(\zeta_{n}-\zeta)\cdot\nabla(\zeta_{n}-\zeta)$ $=a_{\varepsilon}(\Re(\xi_{n});\zeta_{n}-\zeta,\zeta_{n}-\zeta),$
$=a_{\varepsilon}(\Re(\xi_{n});\zeta_{n},\zeta_{n}-\zeta)-a_{\varepsilon}(\Re(\xi_{n});\zeta,\zeta_{n}-\zeta),$
$=l(\zeta_{n}-\zeta)-a_{\varepsilon}(\Re(\xi_{n});\zeta,\zeta_{n}-\zeta),$
$\mathop{\longrightarrow}_{n\to\infty}0-a_{\varepsilon}(\Re(\xi);\zeta,0)=0,$
as $\xi_{n}\to\xi$ strongly in $L^{2}(\Omega;\mathbb{C})$, $\zeta_{n}\rightharpoonup\zeta$ weakly in $V$, and $\zeta_{n}\to\zeta$ strongly in $L^{2}(\Omega;\mathbb{C})$. Take the imaginary part, one has $\zeta_{n}\to\zeta$ strongly in $V$. Also, as the solution of (6) is unique, the full sequence $\left(\zeta_{n}\right)_{n}$ converges towards $\zeta$ in V. We conlude with the Schauder fix point theorem. ∎
Lemma 3.
Let $\delta>0$, $\|\xi_{D}\|_{H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})}\leq\delta$, and $\|g\|_{L^{2}(\partial\Omega_{N};\mathbb{R})}\leq\delta$. For $\delta$ small enough, the solution $\xi_{\varepsilon}$ of (8) is unique.
Proof.
Let $\xi,\hat{\xi}\in V_{D}$. Let $\zeta,\hat{\zeta}\in V_{D}$ be the associated solutions such that for all $\tilde{\zeta}\in V_{0}$,
$a_{\varepsilon}(\Re(\xi);\zeta,\tilde{\zeta})=l(\tilde{\zeta})=a_{\varepsilon}(\Re(\hat{\xi});\hat{\zeta},\tilde{\zeta}).$
Let $\tilde{\zeta}\in V_{0}$. One has
$a_{\varepsilon}(\Re(\xi);\zeta-\hat{\zeta},\tilde{\zeta})$ $=a_{\varepsilon}(\Re(\xi);\zeta,\tilde{\zeta})-a_{\varepsilon}(\Re(\xi);\hat{\zeta},\tilde{\zeta}),$
$=a_{\varepsilon}(\Re(\hat{\xi});\hat{\zeta},\tilde{\zeta})-a_{\varepsilon}(\Re(\xi);\hat{\zeta},\tilde{\zeta}),$
$=\int_{\Omega}\left(\hat{B}(\Re(\hat{\xi}))-\hat{B}(\Re(\xi))\right)\nabla\hat{\zeta}\cdot\nabla\tilde{\zeta}.$
Taking $\tilde{\zeta}=\hat{\zeta}-\zeta$, and considering the imaginary part, one has
$\frac{\varepsilon}{\sqrt{1+C}}\|\zeta-\hat{\zeta}\|_{V}$ $\leq K\||\Re(\hat{\xi})-\Re(\xi)||\nabla\hat{\zeta}|\|_{L^{2}(\Omega;\mathbb{C})},$
$\leq K\|\xi-\hat{\xi}\|_{L^{p}(\Omega;\mathbb{C})}\|\nabla\hat{\zeta}\|_{L^{q}(\Omega;\mathbb{C})},$
$\leq K\|\xi-\hat{\xi}\|_{V}\|\hat{\zeta}\|_{V},$
because $\hat{B}$ is $K$-Lipschitz, and using a generalized Hölder inequality with $\frac{1}{p}+\frac{1}{q}=\frac{1}{2}$, $p>2$, and $q<2$. Using (7) and for $\delta>0$ small enough, one gets that the fix point map is contracting and thus uniqueness. ∎
3 Approximate Discrete problem
This section is dedicated to approximating the solutions of (8). In order to give a convergence proof, we assume in this entire section that there exists a unique solution $\xi_{\varepsilon}$ of (8). This can be ensured by verifying the hypothesis of Lemma 3.
3.1 Discrete setting
Let $\left(\mathcal{T}_{h}\right)_{h}$ be a family of quasi-uniform and shape regular triangulations [10], perfectly fitting $\Omega$. For a cell $c\in\mathcal{T}_{h}$, let $h_{c}:=\mathrm{diam}(c)$ be the diameter of c. Then, we define $h:=\max_{c\in\mathcal{T}_{h}}h_{c}$ as the mesh parameter for a given triangulation $\mathcal{T}_{h}$. Let $V_{h}:=\mathbb{P}^{1}(\mathcal{T}_{h};\mathbb{C})$, the space of affine Lagrange polynomials with complex values. As the solutions we try to compute are only $H^{1}$, we use the Scott–Zhang interpolator written as $\mathcal{I}_{h}$, see [10] for details. As $\xi_{D}\in H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})$, there exists $\zeta_{D}\in H^{1}(\Omega;\mathbb{R})$, $\zeta_{D}=\xi_{D}$ on $\partial\Omega_{D}$. We define the following solution space,
$V_{hD}:=\{\zeta_{h}\in V_{h};\Re(\zeta_{h})=\mathcal{I}_{h}\zeta_{D}\text{ and}\Im(\zeta_{h})=0\text{ on }\partial\Omega_{D}\},$
and its associated homogeneous space
$V_{h0}:=\{\zeta_{h}\in V_{h};\zeta_{h}=0\text{ on }\partial\Omega_{D}\}.$
3.2 Discrete problem
Let $\varepsilon>0$ and $\xi_{h}\in V_{hD}$. Let us first focus on the following linear problem. We search for $\zeta_{h}\in V_{hD}$, such that
$a_{\varepsilon}(\Re(\xi_{h});\zeta_{h},\tilde{\zeta}_{h})=l(\tilde{\zeta}_{h}),\quad\forall\tilde{\zeta}_{h}\in V_{h0}.$ (9)
Lemma 4.
There exists a unique solution $\zeta_{h}\in V_{hD}$ to (9), and one has
$\|\zeta_{h}\|_{V}\leq C^{\prime}C_{tr}\left(1+\frac{M\sqrt{1+C}}{\varepsilon}\right)\|\xi_{D}\|_{H^{\frac{1}{2}}(\partial\Omega_{D};\mathbb{R})}+\frac{\sqrt{1+C}}{\varepsilon}\|g\|_{L^{2}(\partial\Omega_{N};\mathbb{R})},$ (10)
where $C_{tr}>0$ is the constant from the trace theorem, $C>0$ is the Poincaré constant, and $C^{\prime}>0$ is the interpolation constant, see [10].
Proof.
The proof is similar to the proof of Lemma 1 and is omitted for concision. ∎
We now focus on the nonlinear problem consisting of searching for $\xi_{h}\in V_{hD}$,
$a_{\varepsilon}(\Re(\xi_{h});\xi_{h},\tilde{\zeta}_{h})=l(\tilde{\zeta}_{h}),\quad\forall\tilde{\zeta}_{h}\in V_{h0}.$ (11)
Proposition 5.
There exists a solution $\xi_{h}\in V_{hD}$ to (11).
Proof.
Let $\xi_{h}\in V_{hD}$. We define the fix point map $T_{h}:V_{hD}\to V_{hD}$, such that $T_{h}\xi_{h}=\zeta_{h}$, where $\zeta_{h}$ is a solution of (9). Let
[…]
Using Lemma 4, one has $T_{h}B\subset B$.
Let us show that $T_{h}$ is Lipschitz. Let $\xi_{h},\hat{\xi}_{h}\in V_{hD}$ and let $\zeta_{h}:=T_{h}\xi_{h}$, and $\hat{\zeta}_{h}:=T_{h}\hat{\xi}_{h}$. Let $\tilde{\zeta}_{h}\in V_{h0}$, one has
$a_{\varepsilon}(\Re(\xi_{h});\zeta_{h}-\hat{\zeta}_{h},\tilde{\zeta}_{h})$ $=a_{\varepsilon}(\Re(\xi_{h});\zeta_{h},\tilde{\zeta}_{h})-a_{\varepsilon}(\Re(\xi_{h});\hat{\zeta}_{h},\tilde{\zeta}_{h}),$
$=a_{\varepsilon}(\Re(\hat{\xi}_{h});\hat{\zeta}_{h},\tilde{\zeta}_{h})-a_{\varepsilon}(\Re(\xi_{h});\hat{\zeta}_{h},\tilde{\zeta}_{h}),$
$=\int_{\Omega}\left(\hat{B}(\Re(\hat{\xi}_{h}))-\hat{B}(\Re(\xi_{h}))\right)\nabla\hat{\zeta}_{h}\cdot\nabla\tilde{\zeta}_{h}.$
Taking $\tilde{\zeta}_{h}=\hat{\zeta}_{h}-\zeta_{h}$, and considering the imaginary part, one has
$\frac{\varepsilon}{\sqrt{1+C}}\|\zeta_{h}-\hat{\zeta}_{h}\|_{V}$ $\leq K\||\Re(\hat{\xi}_{h})-\Re(\xi_{h})||\nabla\hat{\zeta}_{h}|\|_{L^{2}(\Omega;\mathbb{C})},$
$\leq K\|\xi_{h}-\hat{\xi}_{h}\|_{L^{p}(\Omega;\mathbb{C})}\|\nabla\hat{\zeta}_{h}\|_{L^{q}(\Omega;\mathbb{C})},$
$\leq K\|\xi_{h}-\hat{\xi}_{h}\|_{V}\|\hat{\zeta}_{h}\|_{V},$
because $\hat{B}$ is $K$-Lipschitz, and using a generalized Hölder inequality with $\frac{1}{p}+\frac{1}{q}=\frac{1}{2}$, $p>2$, and $q<2$. Thus $T_{h}$ is continuous and the Brouwer fix point [4] gives the desired result. Regarding uniqueness, $T_{h}$ becomes contracting for $\delta>0$ small enough, which ensures uniqueness. ∎
3.3 Convergence
We study the convergence of the method when $h\to 0$, and for $\varepsilon>0$ fixed. We write the sequence of solutions of (11) as $\left(\xi_{h}\right)_{h>0}$.
Theorem 6.
When $h\to 0$, $\xi_{h}\to\xi_{\varepsilon}$, solution of (8), strongly in $V$.
Proof.
As $\left(\xi_{h}\right)_{h>0}$ is bounded in $V_{D}$, according to Lemma 4, there exists $\zeta_{\varepsilon}\in V_{D}$, up to a subsequence, $\xi_{h}\rightharpoonup\zeta_{\varepsilon}$, weakly in $V$, when $h\to 0$. Using a compact Sobolev injection, one has $\xi_{h}\to\zeta_{\varepsilon}$, strongly in $L^{2}(\Omega;\mathbb{C})$, when $h\to 0$. Let $\tilde{\zeta}\in V_{0}$. Testing (11) with $\mathcal{I}_{h}\tilde{\zeta}\in V_{h0}$, one has
$\int_{\Omega}(\hat{B}(\Re(\xi_{h}))+i\varepsilon)\nabla\xi_{h}\cdot\overline{\nabla\mathcal{I}_{h}\tilde{\zeta}}=a_{\varepsilon}(\Re(\xi_{h});\nabla\xi_{h},\mathcal{I}_{h}\tilde{\zeta})=l(\mathcal{I}_{h}\tilde{\zeta}).$
Thus, as $\hat{B}$ is $K$-Lipschitz and therefore continuous, one has
$\int_{\Omega}(\hat{B}(\Re(\zeta_{\varepsilon}))+i\varepsilon)\nabla\zeta_{\varepsilon}\cdot\overline{\nabla\tilde{\zeta}}=a_{\varepsilon}(\Re(\zeta_{\varepsilon});\nabla\zeta_{\varepsilon},\tilde{\zeta})=l(\tilde{\zeta}).$
And therefore, $\zeta_{\varepsilon}\in V_{D}$ is a solution of (11) which has $\xi_{\varepsilon}\in V_{D}$ as a unique solution. Thus, $\zeta_{\varepsilon}=\xi_{\varepsilon}$, and the full sequence $(\xi_{h})_{h>0}$ converges towards $\xi_{\varepsilon}$. Let us now show the strong convergence of the gradients. As a consequence of the weak convergence, one has
$\liminf_{h\to 0}\|\nabla\xi_{h}\|_{L^{2}(\Omega;\mathbb{C})}\geq\|\nabla\xi_{\varepsilon}\|_{L^{2}(\Omega;\mathbb{C})}.$
Testing (11) with $\hat{\xi}_{h}-\mathcal{I}_{h}\xi_{\varepsilon}\in V_{h0}$, one has
$a_{\varepsilon}(\Re(\xi_{h});\xi_{h},\xi_{h}-\mathcal{I}_{h}\xi_{\varepsilon})=l(\xi_{h}-\mathcal{I}_{h}\xi_{\varepsilon}).$
Therefore,
$a_{\varepsilon}(\Re(\xi_{h});\xi_{h},\xi_{h})$ $=l(\xi_{h}-\mathcal{I}_{h}\xi_{\varepsilon})+a_{\varepsilon}(\Re(\xi_{h});\xi_{h},\mathcal{I}_{h}\xi_{\varepsilon}),$
$\mathop{\longrightarrow}_{h\to 0}0+a_{\varepsilon}(\Re(\xi_{\varepsilon});\xi_{\varepsilon},\xi_{\varepsilon}),$
using the weak and strong convergence towards $\nabla\xi_{\varepsilon}$ and $\xi_{\varepsilon}$. Taking the imaginary part, one has
$\limsup_{h\to 0}\|\nabla\xi_{h}\|^{2}_{L^{2}(\Omega;\mathbb{C})}\leq\|\nabla\xi_{\varepsilon}\|^{2}_{L^{2}(\Omega;\mathbb{C})}$
Consequently, $\nabla\xi_{h}\to\nabla\xi_{\varepsilon}$ strongly in $L^{2}(\Omega;\mathbb{C})$, when $h\to 0$. ∎
4 Numerical tests
dolfinx is used in its configuration for complex numbers, see [26, 25, 13, 1]. The numerical experiments consist in solving (11) for various boundary conditions and different material parameters. (11) is solved through Newton iterations. Then, $\gamma_{h}$, and $y_{\mathrm{eff},h}$ are computed from $\Re(\xi_{h})$ through least-squares and Equations (2) and (1). These results are then compared to experimental results. The regularization parameter $\varepsilon$ is chosen to be as small as possible while still providing a good agreement with the experimental results. The experimental data are recovered using the image recognition capabilities of matlab, see [30, Supplementary Material 7]. The code can be found in [15].
4.1 Auxetic kirigami
The pattern is sketched in Figure 2.
Figure 2: Auxetic kirigami: sketch from [30].
For this pattern, one has $\alpha=-0.9$ and $\beta=0.9$. For $\xi\in[0,\frac{\pi}{3}]$, $-\Gamma_{21}(\xi)>0$ and thus (4) is strictly elliptic. The domain is $\Omega=(0,L)\times(0,L)$, where $L=1.5$, see Figure 3.
$\partial\Omega_{N}$$\partial\Omega_{N}$$\partial\Omega_{D}$$\partial\Omega_{D}$
Figure 3: Auxetic kirigami: geometry of the domain.
Homogeneous Neumann boundary conditions are imposed on the top and bottom surfaces and the nonhomogeneous Dirichlet condition $\xi_{D}$ is imposed on the left and right surfaces.
We consider a mesh of size $h=0.005$, with $119,817$ dofs. The initial guess is chosen as $\xi^{0}_{h}=0$ in $\Omega$. The regularization parameter is chosen as $\varepsilon=0$, as this problem remains elliptic for all $\xi$. $7$ Newton iterations are necessary to reach convergence with a relative tolerance on the residual $r_{tol}=10^{-8}$. Figure 4 shows the numerical results in the deformed configuration against the experimental results.
Figure 4: Auxetic kirigami: left: $\Re(\xi_{h})$, right: experimental result.
The numerical results are very similar to the ones obtained in [30]. One can notice some slight differences with the experimental results. We interpret these as being due to elasticity effects, as described in [31].
4.2 Non-auxetic kirigami
The pattern is sketched in Figure 5.
Figure 5: Non-auxetic kirigami: sketch from [30].
For this pattern, one has $\alpha=-0.9$ and $\beta=0$. For $\xi\in[0,\frac{\pi}{3}]$, $-\Gamma_{21}(\xi)\leq 0$ and thus (4) is degenerate hyperbolic. The domain is $\Omega=(0,L)\times(0,L)$, where $L=1.5$. The Dirichlet boundary conditions are similar to Figure 3.
We consider a mesh of size $h=5.3\cdot 10^{-3}$, with $180,601$ dofs. The regularization parameters is chosen as $\varepsilon=0.5$. $4$ Newton iterations are necessary to reach convergence with a relative tolerance on the residual $r_{tol}=10^{-6}$. Figure 6 shows the comparison between the experimental and numerical results.
Figure 6: Non-auxetic kirigami: left: $\Re(\xi_{h})$, right: experimental result.
With respect to the results of [30], this numerical result shows a better agreement with the experimental data as one can see the depression in the middle of the sample, where $\xi$ is lower than in the surroundings. Also, one sees a similar gradient of $\xi$ on the four edges of the sample.
4.3 Mixed type kirigami
The pattern is sketched in Figure 7.
Figure 7: Mixed kirigami: sketch from Paolo Celli.
For this pattern, one has $\alpha=-1.6$ and $\beta=0.4$. The domain is $\Omega=(0,L)\times(0,L)$, where $L=1.5$. The Dirichlet boundary conditions are similar to Figure 3. A main contribution of this paper is to be able to approximate solutions for this pattern, which was not previously possible, see [30].
We consider a mesh of size $h=5.0\cdot 10^{-3}$, with $180,601$ dofs. The regularization parameter is chosen as $\varepsilon=0.071$. $5$ Newton iterations are necessary to reach convergence with a relative tolerance on the residual $r_{tol}=10^{-6}$. Figure 8 shows the comparison between the experimental and numerical results.
Figure 8: Mixed type kirigami: left: $\Re(\xi_{h})$, right: experimental result.
This numerical test shows a good agreement with the experimental results. The sample seems to deform less than in the experiment which could be due to elasticity effects or indicate a need for more accurate boundary conditions.
5 Conclusion
This paper has presented the analysis of an approximation of a nonlinear degenerate sign-changing divergence form PDE that models the deformation of a specific type of kirigami called rhombi-slits. Under appropriate boundary conditions, it proved existence and uniqueness of solutions to (8). Then, a numerical method based on Lagrange $\mathbb{P}^{1}$ finite elements with complex values is analyzed and shown to converge towards the solutions of (8). Finally, numerical results show the robustness of the method with respect to previous results and experimental results. Future work include trying to solve for an imposed $y_{\mathrm{eff}}$ on the boundary $\partial\Omega_{D}$ and not a $\xi$. Future work could also include working on other kirigami patterns.
Code availability
Acknowledgment
The author would like to thank Paul Plucinsky (University of Southern California) and Ian Tobasco (University of Illinois Chicago) for fruitful discussions.
The author would also like to thank Paolo Celli (Stony Brook University) for providing him with experimental results, produced for [30], and with a matlab code to extract the true deformations from the experiments.
Funding
This work is supported by the US National Science Foundation under grant number OIA-1946231 and the Louisiana Board of Regents for the Louisiana Materials Design Alliance (LAMDA).